### v2

# Interpret the Internal States of Recommendation Model with Sparse Autoencoders

Abstract.
Recommendation model interpretation aims to reveal models’ calculation process, enhancing their transparency, interpretability, and trustworthiness by clarifying the relationships between inputs, model activations, and outputs. However, the complex, often opaque nature of deep learning models complicates interpretation, and most existing methods are tailored to specific model architectures, limiting their generalizability across different types of recommendation models. To address these challenges, we propose RecSAE, an automated and generalizable probing framework that interprets Recommenders with Sparse AutoEncoder. It extracts interpretable latents from the internal states of recommendation models and links them to semantic concepts for interpretation. RecSAE does not alter original models during interpretation and also enables targeted de-biasing to models based on interpreted results. Specifically, RecSAE operates in three steps: First, it probes activations before the prediction layer to capture internal representations. Next, the RecSAE module is trained on these activations with a larger latent space and sparsity constraints, making the RecSAE latents more mono-semantic than the original model activations. Thirdly, RecSAE utilizes a language model to construct concept descriptions with confidence scores based on the relationships between latent activations and recommendation outputs. Experiments on three types of models (general, graph-based, and sequential) with three widely used datasets demonstrate the effectiveness and generalization of RecSAE framework. The interpreted concepts are further validated by human experts, showing strong alignment with human perception. Overall, RecSAE serves as a first step in both model-level interpretations to various types of recommenders without affecting their functions and offering the potential for targeted tuning of models. The code and data are available [^1] https://github.com/Alice1998/RecSAE.
Recommendation Model, Interpretation, Sparse Autoencoder.
ccs: Information systems Recommender systems
1. Introduction
Figure 1. RecSAE employs a probing paradigm to extract interpretable latents from recommenders. The examples are interpretations of SASRec on Amazon.
Transparency is essential for recommender systems as it promotes user trust, fairness, and accuracy (Zhang et al., 2014). Research on recommendation transparency is typically classified into two categories: result-level explanations and model-level interpretations (Zhang et al., 2020a). Result-level approaches focus on explaining relationships between inputs and outputs (Kaffes et al., 2021; Gao et al., 2023), providing users with recommendation explanations. Model-level approaches aim to uncover the decision process of the model itself (Tsang et al., 2020; Du et al., 2023), enabling developers to understand its activations and apply targeted de-biasing to improve model behavior.
While model-level interpretations offer notable advantages, interpreting current recommendation models remains challenging due to the opaque nature of neuron activations in deep learning models (Zhang et al., 2021; Montavon et al., 2018). Most existing model-level interpretations rely on specialized model architectures (Wang et al., 2018b; Hada et al., 2021; Xiao et al., 2017; Du et al., 2023; Zhang et al., 2024), limiting their applicability to different types of recommendation models. As shown in Table 1, three studies aim to develop more generalized interpretation methods. Tsang et al. (Tsang et al., 2020) introduced a feature interaction interpretation method for context-based recommendation models. Another distillation-based approach (Zhang et al., 2020b) learns structural knowledge from an explainable model to embedding-based models. However, it requires joint training with the target model, reducing its adaptability to pre-existing models. Du et al. (Du et al., 2023) propose a clustering-based method, learning interpretable clusters with interaction logs and item tags. However, it sacrifices performance and necessitates model ensemble to compensate. Moreover, the above methods cannot apply targeted modifications to model behaviors. To the best of our knowledge, no existing work provides a generalizable framework for interpreting diverse recommendation models without affecting their performance and simultaneously enabling targeted de-biasing of model behavior.
Table 1. Existing studies on explainable recommendations do not provide model-level interpretations across diverse recommender types and support targeted tuning and de-biasing.
Method Representative Studies Model level General Framework Targeted Tuning
Path Wang 18 (Wang et al., 2018b); Zhao 20 (Zhao et al., 2020) $\sqrt{}$
Textual Hada 21 (Hada et al., 2021); Li 23 (Li et al., 2023) $\sqrt{}$
Attention Xiao 17 (Xiao et al., 2017); Tal 21 (Tal et al., 2021) $\sqrt{}$
Influence Liu 19 (Liu et al., 2019) $\sqrt{}$
Counter- Tan 21 (Tan et al., 2021) $\sqrt{}$
factual Kaffes 21 (Kaffes et al., 2021) $\sqrt{}$
Rule Zhang 24 (Zhang et al., 2024) $\sqrt{}$
RL Wang 18 (Wang et al., 2018a) $\sqrt{}$
LLM Gao 23 (Gao et al., 2023); Lei 23 (Lei et al., 2023) $\sqrt{}$
Distillation Xu 23 (Xu et al., 2023) $\sqrt{}$
Zhang 20 (Zhang et al., 2020b) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$Jointly trained
Feature Tsang 20 (Tsang et al., 2020) $\sqrt{}$ $\sqrt{}\mkern-9.0mu{\smallsetminus}$Context-based
Clustering Du 23 (Du et al., 2023) $\sqrt{}$ $\sqrt{}$
Probing RecSAE (Ours) $\sqrt{}$ $\sqrt{}$ $\sqrt{}$
Motivated by recent advances in employing large sparse autoencoders (SAEs) to interpret transformer-based large language models (Bricken et al., 2023; Templeton et al., 2024; Gao et al., 2024), we utilize SAEs to the interpretation of recommenders. Unlike LLMs, recommendation models exhibit fundamental differences in architecture and behavior, as further discussed in Section 2.2. Various types of recommenders, including general and graph-based models, do not incorporate transformer structures. However, the effectiveness of SAEs for interpretation has been demonstrated primarily within specific structures in transformers.
In this work, we introduce an effective probing paradigm for interpreting various types of recommendation models without altering their functionalities. It also enables target tuning/de-biasing to models. RecSAE operates in three steps: (1) Probe Internal Activations: RecSAE probes the internal activations preceding the recommendation model’s prediction layer to capture its internal representations. (2) Train RecSAE Module: The RecSAE module is trained on these activations with a larger latent space and under sparsity constraints, making RecSAE latents more interpretable than the original model activations. (3) Construct and Verify Semantic Concepts: RecSAE leverages a language model (Llama-3-8b-Instruct (AI@Meta, 2024)) to construct concept descriptions based on input items, latent activations and recommendation model outputs. These concepts are validated by predicting latent activations on new cases and calculating confidence scores based on prediction accuracy. Human experts further verify the constructed concepts and reach an average relevance score of 3.89/4. Additionally, RecSAE enables targeted tuning of the interpreted models by replacing the original activations at the probing position with the reconstructed activations. In summary, our contributions are threefold:
  • We propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. It enables model interpretation without affecting the performance of well-trained models.
  • RecSAE extracts interpretable latents from the internal states of recommendation models and links them to semantic concepts with interpretation confidence scores. These concepts are human-readable and easy to verify.
  • Experiments on three types of recommendation models — general, graph-based, and sequential — across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework. These results highlight RecSAE’s capability to provide model-level interpretations and its potential to facilitate targeted de-biasing of models.
2. Related Work
2.1. Interpretable Recommendation
Table 1 summarizes studies on interpretable and explainable recommendation systems, categorizing methods based on the interpreted subject (result-level or model-level), their generalization (specific to a particular model, a type of model, or across various models), and their ability to enable targeted tuning of the interpreted models.
Result-level interpretation approaches, such as explanation distillation (Xu et al., 2023), LLM augmentation (Gao et al., 2023), and reinforcement learning (Wang et al., 2018a), offer generalizability but do not fully reveal the mechanisms of models. In contrast, model-level interpretation methods generate insights into model structures using techniques like path-reasoning (Wang et al., 2018b; Zhao et al., 2020; Xian et al., 2020), textual modeling (Hada et al., 2021; Li et al., 2021, 2023), attention mechanisms (Xiao et al., 2017; Tal et al., 2021), counterfactual methods (Tan et al., 2021), and clustering (Du et al., 2023). However, these methods are often constrained by dependence on specific model architectures, limiting their generalization. Beyond them, three studies offer a certain degree of generalization. Zhang et al. (Zhang et al., 2020b) transferring structural knowledge from a path-based explainable model. However, as this framework requires target interpretation models to be trained jointly with the explainable model, it limits its adaptability to pre-existing models. Tsang et al. (Tsang et al., 2020) propose a feature-based interpretation method, but the approach is tailored to context-based recommendation models utilizing feature interactions. Du et al. (Du et al., 2023) mine taste clusters from user-item interactions and item tags. However, their clustering paradigm impacts model diversity and performance, requiring ensemble learning with at least three models, thereby increasing training costs.
In summary, existing model-level interpretation methods exhibit limited generalization capabilities to various kinds of recommendation models, and to the best of our knowledge, no related paradigm could utilize interpreted results for targeted model debiasing.
2.2. Neuron Interpretation in Machine Learning
Interpreting deep learning models to understand their decision-making processes is a critical research area (Liang et al., 2021; Jazayeri and Ostojic, 2021; Sajjad et al., 2022; Foote et al., 2023; Fan et al., 2024). The field of neuron interpretation in machine learning has recently advanced through several stages, including interpreting single-layer transformers (Bricken et al., 2023) and analysis of large language models such as GPT-4 and Claude-3 (Gao et al., 2024; Templeton et al., 2024). Specifically, Anthropic has employed dictionary learning (Tošić and Frossard, 2011) to interpret neuron activations in single-layer transformers. Additionally, several studies have analyzed residual stream activations in transformers using sparse autoencoders (Cunningham et al., 2023; Templeton et al., 2024; Gao et al., 2024), aiming to uncover mono-semantic features and improve the interpretability of language models. OpenAI has also leveraged language models to interpret neuron activations at scale (Bills et al., 2023; Gao et al., 2024).
However, recommendation systems exhibit several fundamental differences. Firstly, not all recommendation models are transformer-based. A wide variety of models—such as general, sequential, graph-based, and context-based, content-based models—exist, each with distinct structural characteristics. While sparse autoencoders have demonstrated effectiveness primarily on residual activations within transformer structures (Templeton et al., 2024; Gao et al., 2024), their applicability to other architectures remains underexplored. Additionally, recommendation models are trained on interaction logs to learn collaborative signals, where ID-based training data lack direct correspondence to semantic concepts. Third, the evaluation of interpretability in recommendation systems differs from that in language models, as items are often associated with explicit attributes, and prior work (Du et al., 2023) employed intra- and inter-attribute similarity for assessment. In this paper, we introduce the RecSAE framework, which is designed to interpret various ranges of recommendation models with semantic concepts and provides similarity- and confidence-based scores to quantify the reliability of these interpretations.
Figure 2. RecSAE Framework. It encodes the inner states of recommendation models into large and sparse latent space and constructs corresponding concepts with confidence scores based on the relationship between recommendation input, latent activations, and recommendation outputs. Example recommendation models are LightGCN (He et al., 2020) and SASRec (Kang and McAuley, 2018).
3. RecSAE Framework
3.1. Overview
Formally, the interpretation task is defined as follows. Given 1) a recommendation model, $M_{rec}$; 2) ID-based interaction logs $L=\{(ID_{u},ID_{i}),...\}$ (content-based interaction log are also applicable within the framework); 3) item descriptions $D=\{d_{i},...\}$, such as titles or/and categories, the goal is to interpret what information is encoded in the model. Specifically, we aim to uncover relationships between model internal representations and readable texts.
Based on this general setting, we propose RecSAE, a probing framework that uses a sparse autoencoder structure to interpret the internal states of recommendation models. By encoding model activations into a large, sparse latent space, RecSAE uncovers textual concepts that shape recommendation outcomes, providing insights into previously opaque models. As illustrated in Figure 2, RecSAE features a probing architecture and an automated interpretation pipeline, generating textual descriptions for each concept and validating these interpretations with confidence scores. The following subsection 3.2 and 3.3 provide detailed descriptions.
3.2. RecSAE Training
To achieve interpretable and mono-semantic latents, RecSAE encodes the inner states of recommendation models into a large, sparse latent space through unsupervised learning. This is achieved by using an autoencoder with large latent dimensions and applying sparsity constraints such as the Top-k activation mechanism.
3.2.1. Probing Site
In recommendation models, user representations are typically formed at the final layer before similarity computation with candidate items. This layer captures the user modeling information encoded by the model. Therefore, as shown in Figure 2, we probe at the last layer prior to prediction in this paper. As discussed in Section 6, future work could explore probing at different sites to gain deeper insights. RecSAE is trained to reconstruct the activations at the probing site.
3.2.2. Module Architecture
Unlike standard autoencoders, RecSAE scales the input by a factor of $s$ relative to the original activation dimension and incorporates an additional sparsity objective to promote mono-semanticity (Templeton et al., 2024) in the latent space. A vanilla approach to achieve sparsity is to apply an $L_{1}$ regularization term to the training loss. However, this often results in a dense distribution of low-activation latents, which does not align to achieve $L_{0}$ sparsity in the latent space $S_{Latent}$.
To address this, RecSAE employs a Top-k activation layer, which explicitly controls the number of active latents after encoding. As illustrated by the equations below, $S_{Latent}$ denotes the sparse latent vectors with $k$ latents activated and the others set to 0, and $\hat{S}_{Internal}$ is the reconstructed activations. $M_{Pe}$ and $M_{Pd}$ represent the encoder module and decoder module of the RecSAE module. $W_{enc}$, $b_{pre}$, $W_{dec}$ represent the encoder weight, bias and decoder weight. RecSAE hyper-parameters, scale size $s$ and sparsity factor $k$ are discussed in Section 5.3.
(1) $\begin{split}S_{Latent}&=M_{Pe}(S_{Internal})=\text{Top-k}(W_{enc}(S_{Internal}-b_{pre}))\\ \hat{S}_{Internal}&=M_{Pd}(S_{Latent})=W_{dec}S_{Latent}+b_{pre}\end{split}$
3.2.3. Training Loss
RecSAE adopts an unsupervised learning approach, minimizing the $L_{2}$ distance between $S_{Internal}$ and $\hat{S}_{Internal}$ to train the module. This process ensures that the large, sparse latent space $S_{Latent}$ effectively encodes the information from the internal activations. Thus, interpretations of the latent space reflect information encoded in the interpreted model.
In addition to the primary optimization objective, training RecSAE with Top-k activation introduces a key challenge: preventing ”dead” latents, units that remain inactive during training or inference, to fully utilize the expanded latent space. While activating $k$ latents per input improves efficiency and interpretability, it also means that inactive latents receive no gradient updates from the reconstruction loss and therefore are not effectively trained. To address this, we introduce an auxiliary loss, drawing inspiration from previous work (Gao et al., 2024; Jermyn and Templeton, 2024). The auxiliary loss targets these inactive latents during training, guiding their learning even though they are not involved in the main reconstruction process. Formally, given the reconstruction error $e$, the auxiliary loss $\mathcal{L}_{aux}$ is defined as the $L_{2}$ distance between the main reconstruction error and the reconstruction using top $k_{auxD}$ dead latents, as shown in the equations below. Since the encoder preceding the Top-k activation layer is shared, adding this auxiliary loss incurs limited computational overhead. The training loss is the linear combination of $\mathcal{L}$ and $\mathcal{L}_{aux}$. The effectiveness of RecSAE training designs is validated in Section 5.4.
(2) $\begin{split}\text{loss}&=\mathcal{L}+\alpha\mathcal{L}_{aux}\\ \mathcal{L}&=||S_{Internal}-\hat{S}_{Internal}||^{2}_{2}\\ \mathcal{L}_{aux}&=||e-\hat{e}||^{2}_{2},\quad\text{where }e=S_{Internal}-\hat{S}_{Internal},\\ \hat{e}&=W_{dec}S_{L}^{\prime},\ S_{L}^{\prime}=\text{Top-k}_{auxD}(W_{enc}(S_{Internal}-b_{pre})).\end{split}$
With the above structure and training strategy, RecSAE is capable of probing various types of recommendation models. While applied to well-trained models, RecSAE allows for interpretation without altering model functions. Moreover, by modifying latent activations and replacing the original activations with their modified reconstructions, RecSAE facilitates targeted de-biasing of models. This potential is discussed in Section 5.5.
3.3. RecSAE Concept Construction
After training, RecSAE establishes the mapping relationships between inputs, RecSAE latent activations, and model outputs. Using these links, we design an automated, closed-loop process for generating and validating interpretations for each latent.
RecSAE uses the test interactions for concept construction and verification. As shown in the lower part of Figure  2, it first records the input, latent activations, and recommendation outputs during base model inference on the test set. Next, it divides the continuous, unbounded activation values in the latent space into 10 intensity levels (ranging from 1 to 10), with no activation corresponding to level 0. Then, RecSAE selects the top 2N most activated cases for each latent, randomly assigning N cases for concept construction and reserving the remaining N for verification. Additionally, N non-activated cases for each latent are randomly selected for verification. In line with OpenAI’s setting to interpret neurons with language models (Bills et al., 2023), we set N=5 in this study.
For concept construction, RecSAE generates textual descriptions for each latent, transforming latents into textual concepts, as detailed in Subsection 3.3.1. For concept verification, RecSAE evaluates the interpretations by predicting activation levels on new inputs and calculating precision and recall scores to determine interpretation confidence, as outlined in Subsection 3.3.2.
3.3.1. Concept Description Construction
For latent interpretations, we analyze the highly activated cases. RecSAE incorporates item meta information (title and category if available) to form textual sequences for semantic concept constructions. Each sequence consists of user history interacted items, corresponding model prediction item and its activation level on the interpreted latent.
As shown in Table 5, RecSAE latents effectively extract items with notable similarity. Conseuently, summarizing these similarities into coherent descriptions is a relatively straightforward task. Therefore, instead of relying on paired APIs with state-of-the-art large language models, we opt for a locally deployable model, Llama-3-8B-Instruct (AI@Meta, 2024), to generate concept descriptions in this paper. The instruction starts with a system prompt outlining the latent interpretation task, followed by a one-shot example that includes a query and the corresponding assistant response of concept descriptions. Then, the interpreted latents’ activation cases are provided for interpretation. The complete instruction with one-shot example for each dataset is presented in our anonymous repository.
3.3.2. Concept Description Verification
With the automatically generated concepts, RecSAE further evaluates their quality through confidence scores. To this end, RecSAE uses concept descriptions to predict latent activations on new instances. Specifically, we use two types of test cases: highly activated cases (distinct from those used for concept constructions), and randomly selected cases with no activation. These two sets allow us to compute recall and precision, which together form the basis of the confidence scores.
The verification process begins with a system prompt that outlines the activation-level prediction task, followed by a one-shot example. This example includes a user query containing the concept description, a test case, and the assistant’s predicted activation level. The template is available in the codebase. After predicting latent activation levels, RecSAE computes a confidence score for each concept. A test case is labeled as positive if its predicted activation exceeds 5; otherwise, it is labeled negative. The confidence score is computed as the proportion of correctly classified cases (i.e., true positives and true negatives) among tests, reflecting the accuracy of the concept description in representing latent semantics. In Section 5.2.3, human experts annotate concept quality with confidence scores $\geq$ 0.8 to further validate the effectiveness of RecSAE’s automatic concept construction and validation pipeline.
4. Experimental Setup
We implement the RecSAE framework based on a publicly available recommendation framework, ReChorus (Li et al., 2024), which is recommended by the RecSys Conference [^2] https://github.com/ACMRecSys/recsys-evaluation-frameworks.
4.1. Dataset
We utilize three widely used datasets: (1) Amazon (Leskovec et al., 2007): the Grocery & Gourmet Food subset of a large-scale Amazon Reviews dataset. (2) MovieLens 1M (Harper and Konstan, 2015): a rating data set from the MovieLens website. (3) LastFM (Schedl et al., 2022): listening records of the music platform Last.fm.
We use the built-in data processing in ReChorus: retain positive interactions, 5-core filter, and leave-one-out train, dev, and test split. Table 2 lists the statistics of three datasets. Both datasets contain ID-based interactions and item textual descriptions.
Table 2. Dataset Statistics
Dataset #Users #Items #Interactions
Amazon 14,681 8,713 151,254
MovieLens 6,032 3,125 574,197
LastFM 8,031 27,065 256,699
For SAE training, we infer the interpreted recommendation model over the training set and record their internal activations at the probing position during model inference. With the same approach, we construct the valid and test set for RecSAE.
4.2. Recommendation Models to be Interpreted
The RecSAE framework is capable of interpreting various kinds of recommendation models. For the following experiments, we interpret three types of recommendation models: general, graph-based, and sequential.
  • BPRMF (Rendle et al., 2012): Byesian personalized ranking.
  • LightGCN (He et al., 2020): graph convolution network for rec.
  • SASRec (Kang and McAuley, 2018): self-attentive sequential recommendation.
  • TiMiRec (Wang et al., 2022): target-interest distillation framework for multi-interest sequential recommendation.
4.3. Baseline
RecSAE is a model-level interpretation framework that is not restricted to specific architectures. As shown in Table 1, the only comparable generalized model-level baseline is ECF (Du et al., 2023), a clustering-based method that extracts taste clusters from interactions and item tags. To further strengthen our comparison, we include the distillation method DistillEmb (Zhang et al., 2020b) as an additional baseline, though it is limited to jointly trained models. We also probe internal activations of base models for interpretation comparisons in Table 3. These 3 baselines allow us to evaluate RecSAE’s effectiveness in performance and interpretability across diverse methods.
4.4. Evaluation Metrics
RecSAE aims to interpret recommendation models, requiring that its latent activations reflect the internal states of the interpreted model. Thus, we evaluate both the reconstruction and interpretation performance of RecSAE latents.
For reconstruction, we evaluate the downstream recommendation performance by hit rate and NDCG at positions of 10 and 20 in test-all settings when the original activations are replaced with the reconstructions in the interpreted model, though this replacement is not needed during interpretation.
For interpretation, we refer to related work (Du et al., 2023) that measures the latent utilization through Silhouettes (Rousseeuw, 1987) (a measure of how similar an object is to its own cluster) and the number of interpreted concepts within different confidence score ranges. Specifically, Silhouettes evaluates intra- and inter-concept similarity. The similarity is measured using the cosine distance between item embeddings in the interpreted models. Intra-concept similarity quantifies the average distances between the two most activated items within each concept. Meanwhile, inter-concept similarity measures the average distances between pairs of distinct concepts. We calculate the above measures in concepts with a confidence score of 0.8 or higher. High intra- and lower inter-concept similarity indicate that the interpreted concepts are well-defined and distinct.
4.5. Hyper-parameter
For recommendation models, we use the tuned settings in the ReChorus framework. The embedding size for recommenders is 64, so sparsity $k$ in RecSAE should be less than 64. We perform hyperparameter tuning over the scale size {8, 16, 32}, the sparsity k {8, 16}, learning rate {1e-5, 5e-5, 1e-4} and batch size {4, 8, 16}. The impact of scale size and sparsity k is presented in Section 5.3. For final experiments, we set the scale size to 16 (leading to 1024 latents) and k to 8 for reconstruction and interpretation balance. The detailed hyper-parameter settings for interpreting each model are reported in run.sh within the codebase.
Table 3. RecSAE’s Overall Performance. We evaluate RecSAE’s training with downstream performance when replacing original activations with reconstructions. Note that this replacement is not required during interpretation. For interpretation performance, we measure the item similarity within and between concepts and evaluate interpreted concept numbers (All, confidence score = 1, $\geq$0.9, $\geq$0.8). The best metric on each dataset is bold and the second best is underlined.
Dataset Rec Model Interpret Method Reconstruction Interpretation
Downstream Performance ↑ Similarity #Concepts ↑
[…]
Amazon BPRMF BaseModel 0.0386 0.0200 0.0577 0.0248 0.2362 -0.0670 1 1 5 64
RecSAE 0.0373 0.0192 0.0563 0.0240 0.3056 0.0255 5 16 57 1009
LightGCN BaseModel 0.0426 0.0224 0.0667 0.0284 0.1519 -0.1284 0 0 2 64
RecSAE 0.0413 0.0212 0.0638 0.0268 0.4538 0.0117 8 35 100 1020
SASRec BaseModel 0.0678 0.0367 0.1046 0.0460 0.3483 0.0845 0 5 10 64
RecSAE 0.0624 0.0334 0.0962 0.0419 0.7143 0.0953 46 136 265 1004
TiMiRec BaseModel 0.0649 0.0355 0.0939 0.0427 0.3140 0.0824 0 3 10 64
RecSAE 0.0587 0.0314 0.0885 0.0389 0.6288 0.0529 22 70 187 994
MovieLens BPRMF BaseModel 0.0794 0.0382 0.1386 0.0530 0.1796 0.1075 0 0 2 64
RecSAE 0.0724 0.0355 0.1300 0.0499 0.3795 0.1088 5 18 65 929
LightGCN BaseModel 0.0771 0.0373 0.1343 0.0516 0.4324 0.0742 0 5 8 64
RecSAE 0.0746 0.0373 0.1272 0.0505 0.6225 0.1355 6 36 72 701
SASRec BaseModel 0.2008 0.1049 0.3012 0.1302 0.4445 0.0965 1 1 4 64
RecSAE 0.1868 0.0973 0.2793 0.1205 0.6980 0.1086 8 36 112 970
TiMiRec BaseModel 0.2112 0.1119 0.3105 0.1369 0.5119 0.1143 0 4 7 64
RecSAE 0.2014 0.1083 0.3007 0.1333 0.6564 0.1061 4 31 81 804
LastFM BPRMF BaseModel 0.1139 0.0702 0.1478 0.0787 0.1374 0.0244 0 1 4 64
RecSAE 0.0893 0.0544 0.1175 0.0616 0.3019 0.0162 11 38 112 1019
LightGCN BaseModel 0.1485 0.0952 0.1873 0.1049 0.1959 0.1225 0 0 3 64
RecSAE 0.1203 0.0755 0.1558 0.0844 0.4667 0.0144 7 50 108 1024
SASRec BaseModel 0.2140 0.1579 0.2446 0.1656 0.3688 0.0913 1 2 8 64
RecSAE 0.1722 0.1207 0.2043 0.1288 0.5112 0.0486 30 85 229 1024
TiMiRec BaseModel 0.2147 0.1521 0.2486 0.1606 0.2127 0.0556 0 3 11 64
RecSAE 0.1774 0.1188 0.2103 0.1271 0.5332 0.0420 35 102 222 997
5. Experiments and Results
In experiments, we aim to address the following questions: 1) What is RecSAE’s overall performance in interpreting various types of recommendation models? 2) What concepts does RecSAE interpret from different models? 3) How do RecSAE hyper-parameters influence its reconstruction and interpretation performance? 4) Are RecSAE training designs effective? 5) Can RecSAE perform targeted tuning to models?
5.1. RecSAE Overall Performance (Q1)
Table 3 presents RecSAE’s overall performance in reconstruction and interpretation across four models trained on three datasets.
The recommendation performance remains strong when the original activations are replaced with RecSAE reconstructions in the interpreted model. This confirms that the latent space effectively captures most of the information in the interpreted models at the probing positions. It is important to note that RecSAE does not affect the performances of well-trained models, as reconstructions are not intended to replace the original model activations at the probing positions during interpretation.
RecSAE identifies significantly more valid concepts than various base models. While the four recommendation models exhibit limited interpretability in their activations, RecSAE extracts latent representations that offer fine-grained interpretations with high confidence scores. Intra- and inter-concept similarity evaluations further demonstrate RecSAE’s ability to identify distinct and valid concepts. Specifically, intra-concept similarities consistently surpass those of the base models. RecSAE’s inter-concept similarity remains consistently low, indicating well-separations. While some inter-concept similarities exceed those of the original model, this may be due to base models’ limited set of concepts, with fewer than 10 confidence scores exceeding 0.8. Given the small number of concepts, a low inter-concept similarity is expected.
Table 4. Comparison with Model-level Interpretable Baseline. RecSAE on simple BPRMF yields promising performance. ** represents p¡0.05 significance between best and second best.
Dataset Model H@10 N@10 H@20 N@20
Amazon ECF_single 0.0203 0.0104 0.0341 0.0138
ECF_ensemble 0.0232 0.0115 0.0388 0.0154
DistillEmb 0.0236 0.0120 0.0386 0.0158
RecSAE 0.0373** 0.0192** 0.0563** 0.0240**
MovieLens ECF_single 0.0393 0.0183 0.0675 0.0254
ECF_ensemble 0.0393 0.0202 0.0731 0.0287
DistillEmb 0.0628 0.0285 0.1078 0.0397
RecSAE 0.0724** 0.0355** 0.1300** 0.0499**
LastFM ECF_single 0.0317 0.0277 0.0625 0.0463
ECF_ensemble 0.0741 0.0516 0.0872 0.0549
DistillEmb 0.0940 0.0577 0.1249 0.0655
RecSAE 0.0893 0.0544 0.1175 0.0616
Comparisons with baseline, ECF [^3] https://github.com/zealscott/ECF and DistillEmb [^4] https://github.com/yuan-pku/Distilling-Structured-Knowledge-into-Embeddings-for-Explainable-and-Accurate-Recommendation, are shown in Table 4. ECF’s clustering explanation module is jointly trained with recommendation models, which limits model performance, as seen in the ECF_single results. To mitigate this, ECF ensemble models with different initializations. Based on the ensemble range from the paper, we test both 3-model and 5-model ensembles. The 3-model set-up yields better performance, reported as ECF_ensemble in the table. Despite additional training costs, it still underperforms compared to RecSAE’s reconstruction with BPRMF base. DistillEmb is jointly trained with a path-based model for explanation. While it performs well on the LastFM dataset, the improvements are insignificant (p¿0.19), and RecSAE with the LightGCN base significantly outperforms it. Furthermore, RecSAE does not require this replacement during interpretation. Its probing paradigm does not interfere with well-trained models, offering both performance effectiveness and interpretability.
Table 5. Examples of Interpretable Concepts derived from SASRec model on the Amazon, MovieLens and LastFM dataset. Although some result may seem like item clustering based on titles, it is actually formed by the internal states of SASRec. This pure ID-based model, trained without semantic information, encodes human-understandable concepts within its activations.
Dataset Latent ID Item ID Item Title Activation Interpretation
Amazon 193 4582 Betty Crocker Gluten Free Brownie Mix, 16-Ounce Boxes, … 0.85 gluten-free baked goods, with a particular focus on mixes for baked goods, such as brownies, cakes, cookies, and pancake
4575 Betty Crocker Gluten Free Chocolate Chip Cookie Mix, … 0.83
5882 Bisquick Pancake and Baking Mix, Gluten-Free, … 0.82
271 7644 Happy Baby Happy Munchies Rice Cakes Apple - 1.41 oz 0.76 baby food or snacks, specifically organic and healthy options for infants and toddlers,…
5060 Happy Baby Organic Baby Food 3 Homestyle Meals,… 0.75
5057 Happy Tot Organic Baby Food, Sweet Potato, Apple, Carrot, … 0.74
Movie- Lens 29 1085 Star Trek IV: The Voyage Home (1986) (Action, Adventure, Sci-Fi) 0.57 part of a science fiction franchise, particularly the Star Trek franchise, or have a strong science fiction element
1069 Star Trek: First Contact (1996) (Action, Adventure, Sci-Fi) 0.49
1083 Star Trek: The Wrath of Khan (1982) (Action, Adventure, Sci-Fi) 0.23
617 511 Aladdin (1992) (Animation, Children’s, Comedy, Musical) 0.62 feature child-friendly , such as animation, children’s stories, or musicals, often with a strong emotional
1551 Bambi (1942) (Animation, Children’s) 0.56
1311 Anastasia (1997) (Animation, Children’s, Musical) 0.46
LastFM 260 21101 I’ll Follow The Sun - Remastered 2009, … , by The Beatles 0.65 from the Beatles, with a particular focus on songs that are considered some of their most iconic and well-known
9564 Norwegian Wood (This Bird Has Flown) …, by The Beatles 0.65
4795 Strawberry Fields Forever - Remastered 2009, …, by The Beatles 0.62
804 26791 Traust, in album Futha, by Heilung (…, dark folk, nordic folk) 0.76 strong connections to Nordic culture, particularly folk and folk-inspired music, often with a focus on traditional instruments
9656 Fehu, in album Runaljod – Yggdrasil, by Wardruna (folk, nordic folk) 0.63
9660 Solringen, in album Runaljod – Yggdrasil, by Wardruna (nordic folk) 0.60
5.2. RecSAE Interpreted Concepts (Q2)
5.2.1. Case Study
We provide examples of RecSAE concepts to show the inherent interpretability in the RecSAE latents extracted from the recommendation model. The interpretation cases for SASRec on three datasets are presented in Table 5. Notably, items that activate the same latent exhibit strong semantic similarities. In the Amazon grocery dataset, the concepts range from specific food types to particular ingredients. In MovieLens, the latents span from specific movie series to various genres. In LastFM, the interpreted concepts include different singers, albums and music types. These examples highlight the meaningful concepts captured by RecSAE within the internal states of the interpreted recommendation model.
While some results may seem to reflect item clusters based on their descriptions, these ”clusters” are actually formed by the interpreted model. These purely ID-based models encode human-understandable concepts within their activations after training exclusively on user-item interactions, without any semantic information about items or users. This underscores the remarkable ability of collaborative filtering to capture meaningful information from data. RecSAE, employing an unsupervised training approach, reveals these concepts embedded in previously opaque models.
Figure 3. Concept Survey on BPRMF (left) and SASRec (right) trained on Amazon. RecSAE interprets more diverse concepts in models with stronger recommendation performance.
5.2.2. Concept Survey
The case study above demonstrates that RecSAE successfully extracts highly interpretable latents from recommendation models, and the interpreted concepts are both human-readable and reasonable. To further examine RecSAE’s interpretability, we survey the overall interpretation situation, showing that these concepts provide valuable insights into the model’s capabilities. Figure 3 presents the word cloud for BPRMF and SASRec trained on the Amazon dataset, displaying concepts with a confidence score of 1. Notably, SASRec encodes a wide range of concepts, such as ”bean”, ”hazelnut”, ”vegan”, ”beef”, and ”microwaveable”. In this dataset, the general model BPRMF achieves a rank of 0.0200 in NDCG@10, while SASRec ranks higher at 0.0367. Meanwhile, RecSAE extracts more diverse concepts from SASRec than from BPRMF, directly revealing the differences in models’ encoding abilities.
5.2.3. Human Evaluation on Concept Relevance
To further examine RecSAE’s automatic concept construction and verification pipeline, we conduct human evaluations to assess the relevance of generated concepts with items that activate the latents. Following procedures established in previous work (Lei et al., 2023), three human experts performed four-level relevance annotations (1 irrelevant; 2 relevant; 3 contain similarities among some items; 4 contain key similarities among all items) on randomly selected 100 cases with latent confidence scores greater than 0.8. The agreement between the annotators, measured by Cohen’s kappa, is 0.3565, indicating moderate consistency. The average relevance rating is 3.8900 out of 4, reflecting a high perceived quality of the interpreted concepts. These results validate RecSAE’s automatic interpretation pipeline.
5.3. Reconstruction and Interpretation Frontier with Scale Size $s$ and Sparsity Factor $k$ (Q3)
Figure 4. RecSAE achieves optimal reconstructions when the scale size $s$=16, corresponding to 1024 latents (16* 64 base model dimensions). This trend is consistent for both sparsity factor $k$=8 and $k$=16; in the figure, results are shown for $k$=8.
We conduct a series of experiments to examine the impact of hyper-parameters. Scale size $s$ determines latent space dimensions and sparsity factor $k$ controls the sparsity level of latent activations.
Scale Size $s$. We consider the reconstruction performance when choosing RecSAE scale sizes. We first fix the sparsity factor $k$=8 and vary $s$. As shown in Figure 4, reconstruction performance peaks when $s$=16, corresponding to 1024 latent dimensions. This suggests that 1024 latents are sufficient to encode the 64-dimensional dense internal representations of the investigated recommendation models. This pattern holds for both $k$=8 and $k$=16. Based on these observations, we set $s$=16 in experiments.
Figure 5. Impact of Sparsity Factor k on Reconstruction and Interpretation. Note that RecSAE does not affect model performances as the reconstructions do not need to be placed into original models during interpretations.
Sparsity Factor $k$. $k$ controls the number of active latents in the latent space, impacting both the reconstruction quality and the interpretability of the RecSAE framework. Figure 5 illustrates the trade-off of the two factors. LightGCN is not included in the figure as its pattern is similar to BPRMF. It is important to note that RecSAE does not affect model performance, as the reconstructions do not need to replace the original model activations. Instead, it measures training performance of RecSAE to ensure that the latent space encodes the main information of the interpreted model. As shown in the figure, increasing $k$ typically improves reconstruction performance but results in fewer concepts with high interpretation confidence. As $k$ increases, each item/user is encoded by more latents, leading to a more redundant and less semantically distinct latent space. However, in two cases (BPRMF and TiMiRec on Amazon), the curve peaks at $k=8$, likely because smaller values of $k$ may lead to a relatively small number of activations, causing some latents to be insufficiently explained. In experiments, we set $k$=8 as a balance point between reconstruction quality and interpretability.
5.4. Training Effectiveness (Q4)
We further evaluate the effectiveness of RecSAE’s training by measuring reconstruction performance using MSE and assessing downstream recommendation performance on the test dataset. Table 6 presents the results on the Amazon dataset, with the other two datasets showing similar patterns. The notable reduction in MSE losses and significant improvements in recommendation performance from initialization to after training confirms the effectiveness of RecSAE’s training design. These strong reconstruction performances demonstrate that RecSAE latents effectively capture key information from the original model’s activations, while the reconstruction itself is not required during model interpretation.
Table 6. RecSAE Reconstruction Performance before → after training on Amazon, validating the training effectiveness.
Base Model MSE ↓ HR@10 ↑ NDCG@10 ↑
BPRMF 2.1235 → 0.1331 0.0077 → 0.0724 0.0033 → 0.0371
LightGCN 1.5552 → 0.2178 0.0029 → 0.0413 0.0013 → 0.0212
SASRec 2.0029 → 0.1949 0.0111 → 0.1534 0.0051 → 0.0830
TiMiRec 1.4121 → 0.2656 0.0042 → 0.1576 0.0019 → 0.0870
5.5. Potentials for Targeted Debiasing (Q5)
The RecSAE framework has the potential to make predictable changes to model behavior. Table 7 present case studies showing that decreasing and increasing latent activations leads to corresponding model output behavior. Furthermore, RecSAE may uncover concepts related to popularity bias or unfairness toward users and items. By revising these ”harmful” latents with tuned fairer reconstructions, recommenders can be effectively de-biased.
Table 7. Targeted Tuning. We scale latent value and examine whether the concept appears in top 10 recommended items.
Dataset Latent ID Concept Hit@10
-10× Original +10×
Amazon 288 Coconut 0.0189 0.6604 0.9528
MovieLens 463 Horror 0.0000 0.9213 1.0000
LastFM 67 Taylor Swift 0.0000 0.4079 1.0000
6. Discussion and Future work
Complexity. Assuming the probing site in the original model has a dimension of $|D|$ (typically less than hundreds in recommenders), the space complexity is $O(s|D|^{2})$, where $s$ is the scale size of latent space (in our experiments the optimal $s$ is 16). Training and inference have a time complexity of $O(s|D|^{2})$. As recommendation models typically do not scale with embedding dimensions (Sethi et al., 2022), and the complexity of RecSAE depends only on the scale size and embedding dimensions of the interpreted model, rather than the overall complexity of the base recommender, the RecSAE framework holds the potential for interpreting industry-level recommendationers.
Future Work. 1) Generalize to more types of recommendation models, such as content-based models and industry-level systems; 2) Investigate the impact of different probing positions across model layers to better understand information processing dynamics; 3) Conduct detailed evaluations of RecSAE’s utility for targeted tuning and debiasing to improve model fairness.
7. Conclusion
In this work, we propose the RecSAE framework with a probing paradigm for interpreting various types of recommendation models. It extracts interpretable latents from models’ internal activations and links these latents to semantic concepts, which are further validated to obtain confidence scores. These concepts are human-readable and easy to verify. The RecSAE framework represents the first step toward interpreting models without compromising their performances, while also enabling targeted tuning to models based on the interpretation results. Experiments conducted on three types of recommendation models across widely used datasets demonstrate the effectiveness and generalizability of the RecSAE framework.